# On arithmetic terms expressing the prime-counting function and the n-th prime

### v1

@@ Abstract
We present the first fixed-length elementary closed-form expressions for the prime-counting function, $\pi(n)$, and the $n$-th prime number, $p(n)$. These expressions are represented as arithmetic terms, requiring only a fixed and finite number of elementary arithmetic operations from the set: $\{+,-,\cdot,/,x^{y}\}$.
Mazzanti proved that every Kalmar function can be represented by arithmetic terms. We develop an arithmetic term representing the prime omega function, $\omega(n)$, which counts the number of distinct prime divisors of a positive integer $n$. From this term, we find immediately an arithmetic term for the prime-counting function, $\pi(n)$. We utilize these results, along with a new arithmetic term for binomial coefficients and new prime-related exponential Diophantine equations to construct an arithmetic term for the $n$-th prime number, $p(n)$, thereby providing a constructive solution to a fundamental question in mathematics: Is there an order to the primes?
@@ 1 Introduction
The prime numbers are the most fundamental elements in arithmetic, as every natural number greater than $1$ can be expressed as the product of one or more primes. Primes have been studied for millennia, since their treatment arises naturally from the study of counting and multiplication.
@@ 1.1 The order of the primes
A most natural question is if there exists an order to the prime numbers: A deterministic pattern or structure within the natural numbers that dictates the sequence of primes. Formally, one might seek a function $f:\mathbb{N}\to\mathbb{N}$ defined using only a finite and fixed number of elementary arithmetic operations from the set $\{+,-,\cdot,/,x^{y}\}$, which for a given $n$, returns the $n$-th prime number. Such a function, if it exists, would encapsulate any inherent order or pattern in the primes. Yet constructing such a function has proven to be an extraordinary challenge.
To illustrate why, consider the classical Sieve of Eratosthenes, an algorithm for generating prime numbers. Beginning with the natural numbers greater than $1$, the sieve starts at the first prime $p(1)=2$ and crosses out all its multiples. It then moves to next unmarked number $p(2)=3$, crossing out its multiples. The process of advancing and crossing-out continues indefinitely, leaving only the prime numbers unmarked. This sieve is inherently sequential and iterative, as the determination of the $n$-th prime depends not only on the previous prime, but on all primes that precede it. The sequence of primes is infinite. So then, how could one possibly capture this behavior in a finite and fixed-length formula?
Indeed, the problem of discerning a precise order or structure of the prime numbers has remained entirely out of reach. That is, until relatively recent advancements in mathematical logic and computability unveiled a promising approach.
@@ 1.2 Kalmar functions
where (///) denotes division with flooring and the monus operator ($\dot{-}$) denotes bounded subtraction, defined as: $a\dot{-}b=\max(a-b,0)$ (see [16, pg. 141]). The use of bounded subtraction ensures that outputs remain in $\mathbb{N}$, though standard subtraction (--) may be used equivalently, provided that the output remains a natural number. The $\bmod$ operation is implicitly included in the set, since $a\bmod b=a-b\left\lfloor a/b\right\rfloor$. It is important to emphasize that all arithmetic terms are elementary closed-form expressions and, more restrictively, are of fixed-length, meaning they do not permit variable-length summations and products like those typically allowed in elementary closed-form expressions. Precisely, arithmetic terms are fixed-length elementary closed-form expressions. There is no unanimous consensus of what constitutes a closed-form, however see [19] for an attempt at a rigorous definition.
@@ 1.4 Formulas for primes
Although this formula provides an explicit expression for determining primes, it is not an arithmetic term due to its reliance on summations of variable length. It operates by performing an obfuscated primality test on each number in the sum, based on Wilson’s theorem. The key component, $\cos(\frac{(j-1)!+1}{j}\pi)^{2}$, encodes the $(n-1)!^{2}\bmod\;n$ operation using trigonometric functions, evaluating to $1$ if $j$ is prime and 00 otherwise. This result is then used to locate $p(n)$. There are some well-known variants of Willans formula, such as that of Jones [8]. It is worth mentioning that Jones’ formula is technically an elementary closed-form, though like Willans’ formula, it contains a variable length summation and is therefore also not an arithmetic term. While clever, such formulas rely previous knowledge of the primality of $p(n)$; hence, they are entirely self-referential, and deduce $p(n)$ only by circular means.
@@
$f(n)=2+\left((2\cdot n!)\bmod(n+1)\right)=\begin{cases}n\quad\textup{ if }n\textup{ is prime}\\ 2\quad\textup{ otherwise}\end{cases},$ (2)
@@
Their function does not directly map $n$ to $p(n)$. Instead, the primes $p(n)$ are interspersed with variable-length sequences of $2$s, as determined by the gaps between consecutive primes:
@@ 1.5 Open questions concerning prime formulas
We propose a constructive solution to Question 1 with an arithmetic term for $n$-th prime $p(n)$ that does not rely on any previous knowledge of $p(n)$ nor its primality. From such term, we find immediately a recurrence relation to calculate $p(n+1)$ given $p(n)$, thereby providing a constructive solution to Question 2. The arithmetic terms we find to solve these problems are immense in size and computationally impractical. However, they are composed of a fixed and finite number of elementary arithmetic operations and universally applicable to all $n$. The status of these questions remains unclear, and it is impossible to know for certain what Hardy and Wright meant by their original questions, however we believe that our results provide satisfactory answers and a definitive resolution. As is often the case, our results also raise new important questions, most notably:
@@ 2 Preliminaries
Under $\mathbb{N}$ we understand the set of natural numbers including 00.
@@ 2.1 Number theoretic arithmetic terms
The following number theoretic arithmetic are used by Mazzanti and Marchenkov in [17, 18]:
@@
Here, $\nu_{2}(n)$ represents the $2$-adic order of $n$, which is highest exponent of $2$ dividing $n$. $\operatorname{HW}(n)$ denotes the Hamming weight of $n$, which is the number of $1$s in the binary representation of $n$. $\gcd(a,b)$ is the greatest common divisor of $a$ and $b$. A much simpler arithmetic term for $\gcd(a,b)$ was proposed by Prunescu and Shunia in [13]:
@@
The above arithmetic term for the binomial coefficient $\binom{a}{b}$ is well-known and was originally proved by Julia Robinson in [10]. In § 3.1, we prove two new arithmetic terms for $\binom{a}{b}$ that are of an entirely new construction and greatly simplify our final arithmetic term for $p(n)$.
@@ 2.2 Generalized geometric progressions and the hypercube method
$G_{r}(q,t)=\sum_{j=0}^{t}j^{r}q^{j}.$ (9)
As described by Matiyasevich in the appendix of [21], $G_{r}(q,t)$ can be calculated via the formula
$G_{r+1}(q,t)=\frac{\partial}{\partial q}G_{r}(q,t+1)-\sum_{j=0}^{r}\binom{r+1}{j}G_{j}(q,t).$ (10)
The hypercube method was discovered by Mazzanti [17] and also used in [18, 15].
Let $a,b\in\mathbb{N}:0\leq a<2^{u(\vec{n})}$. We define the function
@@
Now, let $\vec{n}\in\mathbb{N}^{s}$ and consider the integer lattice points contained in the $k$-dimensional cube $[0,t(\vec{n})-1]^{k}$. Define the function
@@
Observe that $\beta(\vec{n},\vec{x})$ provides a bijective mapping between the points in $\{0,\dots,t(\vec{n})-1\}^{k}$ and the integers from 00 to $t(\vec{n})^{k}-1$. In other words, $\beta(\vec{n},\vec{x})$ enumerates the elements of $\{0,\dots,t(\vec{n})-1\}^{k}$ in lexicographical order, assigning values from 00 to $t(\vec{n})^{k}-1$. Let
@@
Therefore, if $W(\vec{n})$ could be expressed as an arithmetic term in $t(\vec{n})$ and $u(\vec{n})$, then the number of zeros of the function $P(\vec{n},\vec{x})$ could also be expressed in such terms. This scenario occurs when $P(\vec{n},\vec{x})$ is an exponential polynomial simple in-$\vec{x}$.
We define a simple monomial in-$\vec{x}$ as an expression of the form
@@
where $r_{1},\ldots,r_{k}\geq 0$, $v_{1},\ldots,v_{k}\geq 1$ are integers, and $c\in\mathbb{Z}$. An exponential polynomial simple in-$\vec{x}$ is defined as a sum of such simple monomials.
@@
$\sum_{\vec{a}\in\{0,\dots,t(\vec{n})-1\}^{k}}a_{1}^{r_{1}}v_{1}^{a_{1}}\cdots a_{k}^{r_{k}}v_{k}^{a_{k}}=G_{r_{1}}(v_{1},t(\vec{n})-1)\cdots G_{r_{k}}(v_{k},t(\vec{n})-1),$
where $G_{r_{i}}(v_{i},t(\vec{n})-1)$ is the sum function corresponding to each variable $a_{i}$. The contribution of a simple exponential monomial $m(\vec{n},\vec{x})$ to $W(\vec{n})$ takes the form:
$\mathcal{A}_{k}(m(\vec{n},\vec{x}),t(\vec{n}),u(\vec{n}))$ $=-(2^{u(\vec{n})}-1)\cdot c\cdot\prod_{i=1}^{k}G_{r_{i}}(2^{2u(\vec{n})t(\vec{n})^{i-1}}v_{i},t(\vec{n})-1)$ (12)
@@
Notice that even for $c=0$, the contribution of the free term is nonzero. Hence, for exponential polynomials simple in-$\vec{x}$, the quantity $W(\vec{n})$ can always be expressed as an arithmetic term in $t(\vec{n})$ and $u(\vec{n})$.
To summarize, given a parameter tuple $\vec{n}=(n_{1},\dots,n_{s})\in\mathbb{N}^{s}$ and a non-negative exponential polynomial function $P(\vec{n},\vec{x})$ simple in-$\vec{x}$, where $\vec{x}=(x_{1},\ldots,x_{k})$, such that $P(\vec{n},\vec{x})$ is defined on the integer lattice points within the $k$-dimensional cube $[0,t(\vec{n})-1]^{k}$ and is strictly bounded by $2^{u(\vec{n})}$, the number of zeros within the cube can be expressed as an arithmetic term $d(\vec{n})$ in $t(\vec{n})$ and $u(\vec{n})$.
@@ 2.3 Sums of squares of multivariate polynomials
2.4 Exponential Diophantine singlefold definitions
@@
A relation $R(\vec{x},y)$ has a singlefold (exponential) Diophantine definition
@@
In the special situation that the relation $R(\vec{x},y)$ is a function $f(\vec{x})=y$, the value of $y$ is uniquely determined by the value of $\vec{x}$ as well. Consequently, in this situation the whole tuple $(\vec{y},y)$ is uniquely determined by the value of the tuple $\vec{x}$. We will deal with several singlefold (exponential) Diophantine definitions and most of them define functions.
Now, we will define various singlefold relations with corresponding lemmas. We omit most of the lemma proofs, since they are quite trivial and follow from their definitions and preceding statements.
We always need the operations quotient and remainder. The singlefold relation $z=\left\lfloor x/y\right\rfloor$ is denoted as
@@
The singlefold relation $z=x\bmod y$ is denoted as
@@
A singlefold relation of the divisibility condition $y\mid x$ will be denoted with
@@
A singlefold relation of the indivisibility condition $y\nmid x$ will be denoted with
@@
The expression $y=\nu_{2}(x)$ means that $y$ is the exponent of $2$ in the prime number decomposition of $x$. We will denote this singlefold relation with
@@
A requirement of the hypercube method (§ 2.2) is that all exponentiations within relations must be powers of a fixed integer base (most commonly $2$) with an exponent that is simple in all variables, meaning that the exponent is a linear function in the displayed unknowns [17, 18]. To illustrate, $2^{2x+3y}$ is a valid exponentiation for a hypercube relation, while $x^{y}$ is invalid. Therefore, we will require a general relation which translates exponentiations of the form $z=x^{y}$ to an equivalent form that is suitable for the hypercube method. The singlefold relation $z=x^{y}$ will be written as
@@
From Mazzanti [17], the general exponentiation can be computed using only powers of $2$ by the formula:
@@
To ensure that our definitions are simple in all variables, meaning that the powers of $2$ are linear functions in the displayed unknowns, we introduce a new variables $y_{1}=x_{1}x_{2}+x_{1}+1$, $y_{2}=y_{1}x_{2}$ as the sum of squares $(y_{1}-x_{1}x_{2}-x_{1}-1)^{2}+(y_{2}-y_{1}x_{2})^{2}=0$. After re-writing the exponents for $2^{(x_{1}x_{2}+x_{1}+1)x_{2}}$ and $2^{x_{1}x_{2}+x_{1}+1}=2^{y_{1}}$ in terms of $y_{1},y_{2}$, we obtain
@@
which defines $x_{3}=x_{1}^{x_{2}}$ singlefold. ∎
3 Singlefold definitions for binomial coefficients
We will need a singlefold exponential Diophantine definition of the relation $z=\binom{x}{y}$ for two important constructions: The singlefold Diophantine definitions for the factorial function and respectively for the Hamming weight of a natural number. For our initial construction, we will use the fact that [10]:

### v2

@@ Abstract
We present the first fixed-length elementary closed-form expressions for the prime-counting function, $\pi(n)$, and the $n$-th prime number, $p(n)$. These expressions are arithmetic terms, requiring only a finite and fixed number of elementary arithmetic operations from the set: addition, subtraction, multiplication, integer division, and exponentiation.
Mazzanti proved that every Kalmar function can be represented as an arithmetic term. We develop an arithmetic term representing the prime omega function, $\omega(n)$, which counts the number of distinct prime divisors of a positive integer $n$. From this term, we find immediately an arithmetic term for the prime-counting function, $\pi(n)$. Combining these results with a new arithmetic term for binomial coefficients and novel prime-related exponential Diophantine equations, we manage to develop an arithmetic term for the $n$-th prime number, $p(n)$, thereby providing a constructive solution to the fundamental question: Is there an order to the primes?
@@ 1 Introduction
The prime numbers are the most fundamental elements in arithmetic, as every natural number greater than 111 can be expressed as the product of one or more primes. Primes have been studied for millennia, since their treatment arises naturally from the study of counting and multiplication.
@@ 1.1 The order of the primes
A most fundamental question is if there exists an order to the prime numbers: A deterministic pattern or structure within the natural numbers that dictates the sequence of primes. Formally, one might seek a function $f:\mathbb{N}\to\mathbb{N}$ defined using only a finite and fixed number of elementary arithmetic operations from the set $\{+,-,\cdot,/,x^{y}\}$, which for a given $n$, returns the $n$-th prime number. Such a function, if it exists, would encapsulate any inherent order or pattern in the primes. Yet constructing such a function has proven to be an extraordinary challenge.
To illustrate why, consider the classical Sieve of Eratosthenes, an algorithm for generating prime numbers. Beginning with the natural numbers greater than 111, the sieve starts at the first prime $p(1)=2$ and crosses out all its multiples. It then moves to next unmarked number $p(2)=3$, crossing out its multiples. The process of advancing and crossing-out continues indefinitely, leaving only the prime numbers unmarked. This sieve is inherently sequential and iterative, as the determination of the $n$-th prime depends not only on the previous prime, but on all primes that precede it. The sequence of primes is infinite. So then, how could one possibly capture this behavior in a finite and fixed-length formula?
Indeed, the problem of discerning a precise order for the prime numbers has remained entirely out of reach. That is, until relatively recent advancements in mathematical logic and computability unveiled a promising approach.
@@ 1.2 Kalmar functions
where the monus operator ($\dot{-}$) denotes bounded subtraction, defined as: $a\dot{-}b=\max(a-b,0)$ (see [16, pg. 141]). The use of bounded subtraction ensures that outputs remain in $\mathbb{N}$, though standard subtraction (--) may be used equivalently, provided that the output remains a natural number. The (///) operator denotes integer division and the $\bmod$ operation is implicitly included in the set, since $a\bmod b=a-b\left\lfloor a/b\right\rfloor$. It is important to emphasize that all arithmetic terms are elementary closed-form expressions and, more restrictively, are of fixed-length, meaning they do not permit variable-length summations and products like those typically allowed in elementary closed-form expressions. Precisely, arithmetic terms are fixed-length elementary closed-form expressions. There is no unanimous consensus of what constitutes a closed-form, however see [19] for an attempt at a rigorous definition.
@@ 1.4 Formulas for primes
Although this formula provides an explicit expression for determining primes, it is not an arithmetic term due to its reliance on summations of variable length. It operates by performing an obfuscated primality test on each number in the sum, based on Wilson’s theorem. The key component, $\cos(\frac{(j-1)!+1}{j}\pi)^{2}$, encodes the $(n-1)!^{2}\bmod\;n$ operation using trigonometric functions, evaluating to 111 if $j$ is prime and 0 otherwise. This result is then used to locate $p(n)$. There are some well-known variants of Willans formula, such as that of Jones [8]. It is worth mentioning that Jones’ formula is technically an elementary closed-form, though like Willans’ formula, it contains a variable length summation and is therefore also not an arithmetic term. While clever, such formulas rely previous knowledge of the primality of $p(n)$; hence, they are entirely self-referential, and deduce $p(n)$ only by circular means.
@@
$f(n)=2+\left((2\cdot n!)\bmod(n+1)\right)=\begin{cases}n+1\quad\textup{ if }n+1\textup{ is prime}\\ 2\quad\textup{ otherwise}\end{cases},$ (2)
@@
Their function does not directly map $n$ to $p(n)$. Instead, the primes $p(n)$ are interspersed with variable-length sequences of 222s, as determined by the gaps between consecutive primes:
@@ 1.5 Open questions concerning prime formulas
We propose a constructive solution to Question 1 with an arithmetic term for $n$-th prime $p(n)$ that does not rely on any previous knowledge of $p(n)$ nor its primality. From such term, we find immediately a recurrence relation to calculate $p(n+1)$ given $p(n)$, thereby providing a constructive solution to Question 2. The arithmetic terms we find to solve these problems are immense in size and computationally impractical. However, they are composed of a finite and fixed number of elementary arithmetic operations and universally applicable to all $n$. The status of these questions remains unclear, and it is impossible to know for certain what Hardy and Wright meant by their original questions, however we believe that our results provide satisfactory answers and a definitive resolution. As is often the case, our results also raise new important questions, most notably:
@@ 2 Preliminaries
Under $\mathbb{N}$ we understand the set of natural numbers including 0.
@@ 2.1 Number theoretic arithmetic terms
The following number theoretic arithmetic terms are used by Mazzanti and Marchenkov in [17, 18]:
@@
Here, $\nu_{2}(n)$ represents the 222-adic order of $n$, which is highest exponent of 222 dividing $n$. $\operatorname{HW}(n)$ denotes the Hamming weight of $n$, which is the number of 111s in the binary representation of $n$. $\gcd(a,b)$ is the greatest common divisor of $a$ and $b$. A much simpler arithmetic term for $\gcd(a,b)$ was proposed by Prunescu and Shunia in [13]:
@@
The above arithmetic term for the binomial coefficient $\binom{a}{b}$ is well-known and was originally proved by Robinson in [10]. In § 3.1, we prove two new arithmetic terms for $\binom{a}{b}$ that are of an entirely new construction and greatly simplify our final arithmetic term for $p(n)$.
@@ 2.2 Generalized geometric progressions and the hypercube method
$G_{r}(q,t)=\sum_{j=0}^{t-1}j^{r}q^{j}.$ (9)
As described by Matiyasevich in the appendix of [21], for all $r>0$, $G_{r}(q,t)$ can be calculated effectively via the recurrence formula:
$G_{r}(q,t)=\frac{\partial}{\partial q}G_{r-1}(q,t+1)-\sum_{j=0}^{r-1}\binom{r}{j}G_{j}(q,t).$ (10)
Every $G_{r}(q,t)$ is an arithmetic term in $q$ and $t$.
The hypercube method was discovered by Mazzanti [17] and is also used in [18, 15].
Consider $a,b\in\mathbb{N}:0\leq a<2^{b}$. We define the function
@@
Let $\vec{n}\in\mathbb{N}^{s}$ and let $u(\vec{n}),t(\vec{n})$ be arithmetic terms. Now, consider the integer lattice points contained in the $k$-dimensional cube $[0,t(\vec{n})-1]^{k}$. Define the function
@@
Observe that $\beta(\vec{n},\vec{x})$ provides a bijective mapping between the points in $\{0,\dots,t(\vec{n})-1\}^{k}$ and the integers from 0 to $t(\vec{n})^{k}-1$. In other words, $\beta(\vec{n},\vec{x})$ enumerates the elements of $\{0,\dots,t(\vec{n})-1\}^{k}$ in lexicographical order, assigning values from 0 to $t(\vec{n})^{k}-1$. Let
@@
[…]
@@
[…]
@@
$\sum_{\vec{a}\in\{0,\dots,t(\vec{n})-1\}^{k}}a_{1}^{r_{1}}v_{1}^{a_{1}}\cdots a_{k}^{r_{k}}v_{k}^{a_{k}}=G_{r_{1}}(v_{1},t(\vec{n}))\cdots G_{r_{k}}(v_{k},t(\vec{n}))=\prod_{i=1}^{k}G_{r_{i}}(v_{i},t(\vec{n})),$
where $G_{r_{i}}(v_{i},t(\vec{n}))$ is the sum function corresponding to each variable $a_{i}$. The contribution of a simple exponential monomial $m(\vec{n},\vec{x})$ to $W(\vec{n})$ takes the form:
$\mathcal{A}_{k}(m(\vec{n},\vec{x}),t(\vec{n}),u(\vec{n}))$ $=-(2^{u(\vec{n})}-1)\cdot c\cdot\prod_{i=1}^{k}G_{r_{i}}(2^{2u(\vec{n})t(\vec{n})^{i-1}}v_{i},t(\vec{n}))$ (12)
@@
[…]
@@ 2.3 Sums of squares of multivariate polynomials
2.4 Exponential Diophantine single-fold definitions
@@
A relation $R(\vec{x},y)$ has a single-fold (exponential) Diophantine definition
@@
In the special situation that the relation $R(\vec{x},y)$ is a function $f(\vec{x})=y$, the value of $y$ is uniquely determined by the value of $\vec{x}$ as well. Consequently, in this situation the whole tuple $(\vec{y},y)$ is uniquely determined by the value of the tuple $\vec{x}$. We will deal with several single-fold (exponential) Diophantine definitions and most of them define functions.
Now, we will define various single-fold relations with corresponding lemmas. We omit most of the lemma proofs, since they are quite trivial and follow from their definitions and preceding statements.
We always need the operations quotient and remainder. The single-fold relation $z=\left\lfloor x/y\right\rfloor$ is denoted as
@@
The single-fold relation $z=x\bmod y$ is denoted as
@@
A single-fold relation of the divisibility condition $y\mid x$ will be denoted with
@@
A single-fold relation of the indivisibility condition $y\nmid x$ will be denoted with
@@
The expression $y=\nu_{2}(x)$ means that $y$ is the exponent of 222 in the prime number decomposition of $x$. We will denote this single-fold relation with
@@
A requirement of the hypercube method (§ 2.2) is that all exponentiations within relations must be powers of a fixed integer base (most commonly 222) with an exponent that is simple in all variables, meaning that the exponent is a linear function in the displayed unknowns [17, 18]. To illustrate, $2^{2x+3y}$ is a valid exponentiation for a hypercube relation, while $x^{y}$ is invalid. Therefore, we will require a general relation which translates exponentiations of the form $z=x^{y}$ to an equivalent form that is suitable for the hypercube method. The single-fold relation $z=x^{y}$ will be written as
@@
From Mazzanti [17], the general exponentiation can be computed using only powers of 222 by the formula:
@@
To ensure that our definitions are simple in all variables, meaning that the powers of 222 are linear functions in the displayed unknowns, we introduce a new variables $y_{1}=x_{1}x_{2}+x_{1}+1$, $y_{2}=y_{1}x_{2}$ as the sum of squares $(y_{1}-x_{1}x_{2}-x_{1}-1)^{2}+(y_{2}-y_{1}x_{2})^{2}=0$. After re-writing the exponents for $2^{(x_{1}x_{2}+x_{1}+1)x_{2}}$ and $2^{x_{1}x_{2}+x_{1}+1}=2^{y_{1}}$ in terms of $y_{1},y_{2}$, we obtain
@@
which defines $x_{3}=x_{1}^{x_{2}}$ single-fold. ∎
3 Single-fold definitions for binomial coefficients
We will need a single-fold exponential Diophantine definition of the relation $z=\binom{x}{y}$ for two important constructions: The single-fold Diophantine definitions for the factorial function and respectively for the Hamming weight of a natural number. For our initial construction, we will use the fact that [10]: